What is the molecular formula and structure? Use the mass, infrared, proton NMR, and carbon NMR. I have a structure and formula completed, but I wanted to confirm. TRANSMETTANCE 41 50 4000 3000 2000 DOST 1000 HAVENUHBERI-1Mass of molecular ion: 116 100 80- Relative Intensity 40 20- TI Other 10 20 ITHTHIS 30 40 fittihristutt tartott 50 60 70 80 m200 180 160 140 TTTTTT 120 100 80 60 40 20 ppmppm
Answer #1

IR spectra : between 1700-1760 cm-1 there seems to be two overlapped strong peaks corresponding to carbonyl ( -C=O) group , one is keto group and other might be an ester. - 3000-2900cm-1 peak due to sp3C-H stretching.

Mass spectra : Molar mass : 116 m/z ; base peak at m/z : 43 ==> using IR we acn predict it is due to CH3-C\equivO+: species.

using IR and MS : Molecular formula : C5H8O3.

1H-NMR : relative integration ratio : 3:2:3   ==> 2.2 ppm ( singlet ,3H ) , 3.4 ppm ( S, 2H ) and 3.7 ppm ( S, 3H). ( peak at 4.9 ppm and others might be due to keto-enol tautomerism )

13C-NMR :   total 5 non-equivalnet C : 202 ppm ( keto - ) , 168 ppm (ester carbonyl) , 30,49 and 52 ppm ( sp3 C)

Most probable Molecular structure :      CH3-(C=O)-CH2-(C=O)-OCH3. Methyl acetoacetate.